Use the functions to answer the question.
f(x) = 4x-7; g(x) = x + 3
What is the value of (f°g)(4)?

Answer by rothauserc(4718)

first compute fog = (4x-7) * (x+3) = 4x^2 + 5x - 21
then we compute fog (4) = 4*16 + 20 - 21 = 63